OpenStudy (anonymous):

A number added to one-sixth of itself is 28. Solve and show all working out.

OpenStudy (anonymous):

Ok dokey :D That's easy enough, I already know the answer! Since you are looking for a number that's been multiplied by the function (7/6). That's the same thing as adding 1/6 of the total value. In example, 1 multiplied by (7/6) would equal 7/6. 1/6 added to 1 would also result in the same equation... You can solve for the original input by dividing by 7/6. Y= (7/6)x 28 = (7/6)x 28/(7/6) = x x + 24

OpenStudy (anonymous):

x=24***'

OpenStudy (anonymous):

Sorry, I don't understand where the (7/6) comes from.

OpenStudy (texaschic101):

A number added to 1/6 of itself is 28 Lets x be the number.. x + 1/6x = 28 (you have to have a common denominator to add fractions....that x is the same as 1/1x....so the common denominator is 6, thus changing the 1/1x to 6/6x, then add it to 1/6x and get 7/6) 6/6x + 1/6x = 28 7/6x = 28 x = 28/(7/6) x = 28 * 6/7 x = 168/7 x = 24 check.. x + 1/6x = 28 24 + 1/6(24) = 28 24 + 24/6 = 28 24 + 4 = 28 28 = 28 (correct)

OpenStudy (anonymous):

Thank you!